OpenStudy (anonymous):

The equation 5/x = 2/x + 3/x is true for all values of x where x is not equal to 0. Is this equation an identity? Explain your answer. Can you help me @esshotwired ?

OpenStudy (esshotwired):

is this the equation? \[\frac{ 5 }{ x }=\frac{ 2 }{ x }+\frac{ 3 }{ x }\]

OpenStudy (anonymous):

yes

OpenStudy (esshotwired):

it is true that any value of x greater than 0 will make it true because if you solve this before putting in the value of x it would be \[\frac{ 5 }{ x }=\frac{ 5 }{ x }\]

OpenStudy (esshotwired):

it is an identity http://www.mathsisfun.com/definitions/identity.html

OpenStudy (anonymous):

ok, so it is Identity?

OpenStudy (esshotwired):

because no matter what the variable equals, the solution will always be true

OpenStudy (esshotwired):

yes

OpenStudy (anonymous):

Ok, great, thank you so much for helping me!!!

OpenStudy (esshotwired):

your welcome

OpenStudy (anonymous):

Have a great day/evening/night!!! Bye!!

OpenStudy (esshotwired):

haha you too! btw its afternoon by me ;)

OpenStudy (anonymous):

where do you live?

OpenStudy (anonymous):

state-wise!

OpenStudy (esshotwired):

georgia

OpenStudy (anonymous):

HA, You're right below me, I live in SC!!

OpenStudy (anonymous):

I think you've told me that before, I just forgot!

OpenStudy (esshotwired):

oh haha! cool. what part?

OpenStudy (anonymous):

Columbia!

OpenStudy (anonymous):

you?

OpenStudy (esshotwired):

canton

OpenStudy (anonymous):

Where is that? I've never heard of it!

OpenStudy (esshotwired):

in cherokee county, the county above the county that Atlanta is in

OpenStudy (anonymous):

oh, ok, so you're close to Augusta?

OpenStudy (esshotwired):

not really. give me a sec so i can find a map

OpenStudy (anonymous):

ok! :)

OpenStudy (anonymous):

oh, ok!

OpenStudy (anonymous):

Well, I've gtg, so, bye!! Have a great AFTERNOON!!! :D

OpenStudy (esshotwired):

that is where i am compared to you

OpenStudy (esshotwired):

lol you too!